# Random Schreier graphs as expanders

### v6

@@ Abstract
Expander graphs, due to their good mixing properties, are useful in many algorithms and combinatorial constructions. One can produce an expander graph with high probability by taking a random graph. For example, for the case of bipartite graphs of degree $d$ and $n$ vertices in each part we may take independently $d$ permutations of an $n$-element set and use them for edges.
This construction is much simpler than all known explicit constructions of expanders and gives graphs with good mixing properties (small second largest eigenvalue) with high probability. However, from the practical viewpoint, it uses too many random bits, so it is difficult to generate and store these bits for reasonably large graphs. The natural idea is to replace the group of all permutations by its small subgroup. Let $n$ be $q^{k}-1$ for some $k$ and some prime $q$. Then we may interpret vertices as non-zero $k$-dimensional vector over the field $\mathbb{F}_{q}$, and take random linear permutations, i.e., random elements of $GL_{k}(\mathbb{F}_{q})$. In this way the number of random bits used will be polynomial in $k$ (i.e., the logarithm of the graph size, instead of the graph size itself) and the degree.
In this paper we provide some experimental data that show that indeed this replacement does not change much the mixing properties (the second eigenvalue) of the random graph that we obtain. These data are provided for several types of graphs (undirected regular and biregular bipartite graphs). We also prove some upper bounds for the second eigenvalue (though it is quite weak compared with the experimental results).
Finally, we discuss the possibility to decrease the number of random bits further by using Toeplitz matrices; our experiments show that this change makes the mixing properties of graphs only marginally worse, while the number of random bits decreases significantly.
Source Code:
@@ 1 Introduction
Adding one step to our walk means multiplying the probability distribution by the (normalized) adjacency matrix of the graph. The largest eigenvalue of this matrix corresponds to the stationary state (the uniform distribution), so the convergence speed is determined by the second largest eigenvalue of this matrix. Graphs having a good mixing property are called expanders. They are used in numerous applications in computer science and in coding theory. In these applications we usually need an instance of an expander graph.
How can we get a graph with good mixing properties, i.e., with small second eigenvalue (an expander)? The natural idea is to try a random graph. For example, we may consider $d/2$ random permutations of $n$ vertices (where $d$ is some even number) and use them as edges connecting the vertices before and after the permutation. Since we consider undirected graphs, we get a graph of degree $d$ (that might have parallel edges and loops). This contruction is called the permutation model. The largest eigenvalue equals $d$ (before the normalization that divides all elements by $d$), and all eigenvalues are real (since the matrix is symmetric). What can be said about the distribution of the second largest eigenvalue for a graph randomly sampled in this model?
@@
Observe also that the second largest eigenvalue for some particular graph can be computed efficiently, while a direct computation of the expansion ratio (as well as other combinatorial parameters characterizing the properties of expansion and connectivity) would require an exponential search over all subsets of the graph’s vertices. At the same time, we should notice that the spectral technique has its limitations, and computing the second largest eigenvalue cannot give better guarantee than an expansion ratio greater than half of the degree ([13]).
There are many different explicit constructions of expanders (see, e.g., [16, 9, 2, 19, 21, 18]); some of them also achieve a second eigenvalue close to the optimal. However, they are quite complicated, may have some restrictions on the degree, and the bounds for time complexity and eigenvalues can be of asymptotic nature. The advantage of the random graphs is the simplicity: we generate a graph using a simple randomized algorithm, check its second eigenvalue, and if it is small enough, we can use the graph when good expansion properties are needed. Note that the second eigenvalue needs to be computed only once (and this is still feasible for rather large graphs using modern algorithms and computers), and after that we may safely use the graph in our application.
@@
In this paper we experiment with one approach of this type. Namely, we replace random permutations by linear random permutations. Let $\mathbb{F}_{q}$ be a finite field with $q$ elements. The $k$-dimensional vector space over this field contains $q^{k}$ elements ($k$-tuples of elements of $\mathbb{F}_{q}$). A $k\times k$ invertible matrix determines a permutation of these elements. The zero tuple is a fixed point, so we delete it and get a $q^{k}-1$-element set where the group $GL_{k}(\mathbb{F}_{q})$ of invertible $k\times k$ matrices over $\mathbb{F}_{q}$ acts transitively. The non-zero tuples will be the vertices of the graph, and $d/2$ invertible matrices give us $d/2$ permutations of vertices, so we may construct a regular graph with $q^{k}-1$ vertices and degree $d$.
The construction that we described is a special (and easy to implement) case of a Schreier graph. Let us recall the general definition.
Definition 1.1.
Let $H$ be a group acting transitively on a set $X$. We denote $h.x\in X$ the product of the action of $h\in H$ on $x\in X$. Let $S$ be a random multiset of elements of $H$ of $d$ elements. Then the Schreier (undirected) graph
$G=Sch(H\circlearrowleft X,S)$
is a $2d$-regular graph of size $|X|$ whose edges are $(x,s.x)$ for $x\in X$ and $s\in S$.
The graphs in our construction are Schreier graphs $Sch(H\circlearrowleft X,S)$ where $X=(\mathbb{F}_{q}^{k})^{*}$, $H=GL_{k}(\mathbb{F}_{q})$ acting on $X$ by the usual matrix-vector multiplication, and $S$ is a randomly chosen multiset of group elements.
@@
This approach uses much less bits than the general permutation model: for the case $q=2$ we need about $k^{2}$ bits for each random matrix instead of $n\log n$ bits for a random permutation (and $n$ is about $2^{k}$). Increasing the size of the field, we get additional economy. The problem is that there are no theoretical guarantees for mixing properties: results proven for the random permutation model do not have yet their counterparts for random linear permutations (Though some weaker results exist; see [8], [22] or [10]). Note, however, that we can use this approach in practice safely, assuming we can check the eigenvalues before using the graph.
@@
In this paper we provide some experimental evidence showing that this approach may work for several settings. In Section 2 we compare the empirical distributions for the second eigenvalue for random permutations and linear random permutations. We also compare different choices for the field size. In Section 3 we provide a similar comparison for bipartite graphs including the case where the sizes of the left and right parts differ significantly. To decrease the number of random bits further, we can use Toeplitz matrices; the experiments with them are reported in Section 4. Finally, in Section 5 we prove some upper bounds for the expected second largest eigenvalue of graphs considered in Section 2 and 3. Unfortunately, these bounds are quite complicated (the full proofs are given in the appendix) and quite weak; they are based on tools from [6] that give only $O(d^{3/4})$ bound. Still it can be considered as an additional motivation to consider only linear permutations instead of all permutations.
For reproducibility purposes, we provide the source code link in the front page.
@@ 2 Regular graphs
We consider a finite field $\mathbb{F}_{q}$ with $q$ elements, and the vector space of $k$-tuples $\mathbb{F}_{q}^{k}$ over $\mathbb{F}_{q}$. Then we choose randomly $d/2$ invertible matrices $T_{1},\ldots,T_{d/2}$, and use the corresponding permutations of the set of non-zero tuples to get an undirected graph of degree $d$. More precisely, a non-zero tuple $x\in\mathbb{F}_{q}^{k}$ has neighbors
$T_{1}x,\ldots,T_{d/2}x,T_{1}^{-1}x,\ldots,T_{d/2}^{-1}x.$
This is a regular graph with $q^{k}-1$ vertices and degree $d$ (that may contain loops and parallel edges), and we compute numerically the second eigenvalue of this graph.
This procedure is repeated many times, and then the empirical distribution of the second eigenvalues is shown (Figure 1). For comparison, we show also the distribution for the random permutation model, and for different field size.
Figure 1: To make the figure easier to read, all the eigenvalues measured have been divided by $2\sqrt{d-1}$. The two upper curves (quite close to each other) show the empirical distributions for random permutation model (dashed line) and random linear permutation model for graphs of degree $d=30$ ($15$ permutations) and size $116383=2^{14}-$ 1. The third curve (wider green one in the bottom part) shows the distribution for random linear permutations on $5$-tuples from a $7$-element field, so the graph has $168067^{5}-1=$ 16806 vertices (and the same degree $30$, obtained with $15$ permutations). For each distribution $5000$ experiments were made. The eigenvalues are grouped in $40$ bins
The invertible matrices were generated by choosing a random [^1] The random number generator used is Mersenne twister ([17]); we tested physical random generators without noticing any difference. matrix and then testing whether it is invertible. For Boolean matrices, the fraction of invertible ones is above $1/4$ (see, e.g., [1]), so we do not need too many trials; for bigger fields, the fraction is even bigger. Computing the determinant (checking the invertibility) is easy. For computing the eigenvalues we used the C++ library Spectra [^2] Spectra’s home page: https://spectralib.org/. that implements the Lanczos algorithm [7].
@@
We also tested another set of parameters: we took $5$-tuples from a $7$-element field (so the resulting graph size is close to the previously considered one, $168067^{5}-1=$ 16806 instead of $163832^{14}-1=$ 16383, which makes the results roughly comparable). Note that this choice allows us to decrease the number of random bits used for the graph generation. We see that this is achieved for a price: the distribution is now less concentrated, though still very close to the Ramanujan threshold. (This is not a complete surprise that more random bits lead to better concentration.) One can also observe (we do not have any explanation for this paradoxical observation) that the probability to get a random graph that is on the left of, say, $0.998$, becomes bigger for a bigger field. So, if we try to get one graph with small eigenvalue by making many trials, the bigger field size may be preferable. One could try to go to the extreme in this direction and consider $1\times 1$ invertible matrices, i.e., consider multiplication by non-zero elements in some field $\mathbb{F}_{q}$ (the graph size is then $q-1$). This is, however, a bad idea since in this case the invertible matrices commute and therefore many paths in a random walk arrive at the same place just due to commutativity. Such a graph will have poor mixing properties.
Additionally, we can observe that variations on the degree of the graphs (for a given construction) does not seem to have much impact on the distribution of the second largest eigenvalues. We can observe this on Figure 2. Though the difference is hardly visible, we can see that the typical $10$-regular graph has slightly greater second largest eigenvalue compare to the unit we used ($2\sqrt{d-1}$). This is not surprising since low degree random graphs tend to be worse expanders. However, the eigenvalues are slightly more concentrated.
Figure 2: Second largest eigenvalue distribution of Schreier graphs of $GL_{14}(\mathbb{F}_{2})$ of degree $10,30,60$ and $120$.
@@ 3 Bipartite biregular graphs
We have discussed undirected regular graphs with good mixing properties, but in some cases one may wish to construct a bipartite graph that is biregular (degrees of the vertices are the same in each part, but may differ in left and right parts) and has good mixing properties (for the back and forth random walk).
@@
However, in the bipartite case we do not need to use permutation together with its inverse. The random permutation model thus can be modified: we consider $d$ random permutations instead; each one determines a one-to-one correspondence between parts and we draw edges that represent these $d$ correspondences. We use this construction in our experiments with bipartite graphs (the previous construction would have given the same result as for regular non bipartite graphs). This allows to have bipartite graphs of any (possible odd) degree.
@@
These constructions can be performed both with random permutation and random linear permutations. The same question arise: what can we say about mixing properties of graphs in these two settings? A traditional way to measure mixing property of a bipartite graph is to consider the (non-normalized) adjacency matrix for it (height and width are equal to the total number of vertices in both parts). The eigenvalues of this matrix are grouped into pairs, starting with $\sqrt{d_{L}d_{R}}$, where $d_{L}$ and $d_{R}$ are left and right degrees (the eigenvector is $\sqrt{d_{L}}$ in the left part and $\sqrt{d_{R}}$ in the right part) and $-\sqrt{d_{L}d_{R}}$ (to change the sign of the eigenvalue we change signs in all eigenvector coordinates in one of the parts). The next pair, $\lambda>0$ and $-\lambda<0$, determines the mixing properties, and we call $\lambda$ the second largest eigenvalue (ignoring the negative ones).
@@
Figure 3: Here, n1 is the size of each partition. The dashed line (blue) shows the distribution of second eigenvalues for a bipartite graph constructed from $30$ random permutations of $16383$-element set (nonzero bit vectors of size $14$). The orange line (close to the first one) shows the same for $30$ linear permutations (over $\mathbb{F}_{2}$). The third line (green, in the bottom part) shows the distribution for $30$ permutations of a $16806$-element set randomly chosen among the $\mathbb{F}_{7}$-linear permutations of nonzero $5$-tuples with elements in $\mathbb{F}_{7}$.
@@
For the case of different left and right degrees the results of our numerical experiments are shown in Figure 4.
Comparing this figure with the previous one, we observe some differences (that we cannot explain): the random permutations now give slightly (but visibly) better results compared to random linear permutations. On the other hand, changing the size of the field (and matrix dimension) now has much less effect. In fact, results may depend on which triples we are merging, for the linear case. The choice of the merged triplets of vertices could have some influence in the linear permutation case. This might be due to the additional structure induced by the the construction (compared to general permutations). However, this phenomenon remains to be explained.

### v7

@@ Abstract
Expander graphs, due to their mixing properties, are useful in many algorithms and combinatorial constructions. One can produce an expander graph with high probability by taking a random graph (e.g., the union of $d$ random bijections for a bipartite graph of degree $d$). This construction is much simpler than all known explicit constructions of expanders and gives graphs with good mixing properties (small second largest eigenvalue) with high probability. However, from the practical viewpoint, it uses too many random bits, so it is difficult to generate and store these bits for large graphs. The natural idea is to restrict the class of the bijections that we use. For example, if both sides are linear spaces $\mathbb{F}_{q}^{k}$ over a finite field $\mathbb{F}_{q}$, we may consider only linear bijections, making the number of random bits polynomial in $k$ (and not $q^{k}$).
In this paper we provide some experimental data that shows that this approach conserves the mixing properties (the second eigenvalue) for several types of graphs (undirected regular and biregular bipartite graphs). We also prove some upper bounds for the second eigenvalue (though they are quite weak compared with the experimental results).
Finally, we discuss the possibility to decrease the number of random bits further by using Toeplitz matrices; our experiments show that this change makes the mixing properties only marginally worse while the number of random bits decreases significantly.
@@ 1 Introduction
Adding one step to our walk means multiplying the probability distribution by the (normalized) adjacency matrix of the graph. The largest eigenvalue of this matrix corresponds to the stationary state (the uniform distribution), so the convergence speed is determined by the second largest eigenvalue of this matrix. Graphs having good mixing properties are called expanders. They are used in numerous applications in computer science and in coding theory (see e.g. [13, 24, 23, 9]).
How can we get a graph with good mixing properties, i.e., with small second eigenvalue (an expander)? The natural idea is to try a random graph. For example, for undirected graphs we may consider $d/2$ random permutations of $n$ vertices (where $d$ is some even number) and use them as edges connecting the vertices before and after the permutation. We get a graph of degree $d$ (that might have parallel edges and loops). This construction is called the permutation model. The largest eigenvalue equals $d$ (before the normalization that divides all elements by $d$), and all eigenvalues are real (since the matrix is symmetric). What can be said about the distribution of the second largest eigenvalue for a graph randomly sampled in this model?
@@
There are many different explicit constructions of expanders (see, for example, [16, 10, 2, 19, 21, 18]); some of them also achieve a second eigenvalue close to the optimal. However, they are quite complicated, may have some restrictions on the degree, and the bounds for time complexity and eigenvalues can be of asymptotic nature. The advantage of the random graphs is the simplicity: we generate a graph using a simple randomized algorithm, check its second eigenvalue, and if it is small enough, we can use the graph when good expansion properties are needed. Note that the second eigenvalue needs to be computed only once (and this is still feasible for rather large graphs using modern algorithms and computers), and after that we may safely use the graph in our application.
@@
In this paper we experiment with one approach of this type. Namely, we replace random permutations by linear random permutations. Let $\mathbb{F}_{q}$ be a finite field with $q$ elements. The $k$-dimensional vector space over this field contains $q^{k}$ elements ($k$-tuples of elements of $\mathbb{F}_{q}$). A $k\times k$ invertible matrix determines a permutation of these elements. The zero tuple is a fixed point, so we delete it and get a $q^{k}-1$-element set where the group $\textit{GL}_{k}(\mathbb{F}_{q})$ of invertible $k\times k$ matrices over $\mathbb{F}_{q}$ acts transitively. The non-zero tuples will be the vertices of the graph, and $d/2$ invertible matrices give us $d/2$ permutations of vertices, so we may construct a regular graph with $q^{k}-1$ vertices and degree $d$.
The construction that we described is a special (and easy to implement) case of a Schreier graph: let $H$ be a group acting transitively on a set $X$, so that $h\in H$ maps $x\in X$ into $h.x$. Let $S$ be a random multiset containing $d$ elements of $H$. Then the Schreier (undirected) graph $G$ is a $2d$-regular graph of size $|X|$ whose edges are $(x,s.x)$ for $x\in X$ and $s\in S$. In our example $X=(\mathbb{F}_{q}^{k})^{*}$ and $H=\textit{GL}_{k}(\mathbb{F}_{q})$.
@@
This approach uses much less bits than the general permutation model: we need $O(\log^{2}n)$ bits for each random matrix instead of $O(n\log n)$ bits for a random permutation. The problem is that there are no theoretical guarantees for mixing properties: results proven for the random permutation model do not have yet their counterparts for random linear permutations (though some weaker results exist; see Theorem 1.3 in [11]; see also [22] or [8]). Note, however, that we can use this approach in practice safely, assuming we can check the eigenvalues before using the graph.
How to choose the size of graphs for numerical experiments? On the one hand, experiments with “relatively small” graphs are less predictable (therefore, more interesting), since their spectral and combinatorial properties may be far from the asymptotic limits. On the other hand, for “relatively large” graphs, it is easier to justify the advantage of pseudo-randomness (sampling a pseudo-random graph requires significantly less random bits than a truly random one). In this paper, due to the limited space, we focus on expanders of sizes that can be in demand in practice (we only produced very few large graphs). In fact, the size of a “useful” expander depends on the area of application. One of the most popular field of applications of expander graphs is coding theory. In what follows we mostly focus on graphs with $\approx 16000$ vertices, which could be used to construct expander codes (see [23, 24]) with parameters comparable with the codes used in real life (like, e.g., the “short” variant of the LDPC from the DVB-S2 standard). However, we emphasise the importance of studying expander graphs of other sizes. Indeed, in some practically used codes, the codewords are significantly longer (many tens of thousands of bits). Moreover, we have theoretical evidences that in some settings codes with codewords of size $10^{7}$ and even more can be useful (see, e.g., [9]). This means that we eventually may need practical constructions of expander graphs with dozens of millions of vertices. In Appendix we also mention practically useful graphs with even larger (exponentially larger!) number of vertices.
@@
In this paper we provide some experimental evidence showing that this approach may work for several settings. In Section 2 we compare the empirical distributions for the second eigenvalue for random permutations and linear random permutations for undirected graphs. We also compare different choices for the field size. In Section 3 we provide a similar comparison for bipartite graphs including the case where the sizes of the left and right parts differ significantly. To decrease the number of random bits further, we can use Toeplitz matrices; the experiments with them are reported in Section 4. Finally, in Section 5 we show how the tools from [6] can work for random linear permutation and sketch the proof of some (weak) upper bound for the expected second largest eigenvalue of Schreier graphs (of any group) and for graphs from Toeplitz matrices. With some additional work, these tools provide better bounds (though still quite weak compared with the experimental results) for graphs considered in section 2 and 3 of a given size, and we discuss these bound. Unfortunately, even the exact statements of these bounds are quite complicated and the details (as well as the proofs) are given in Appendix.
@@ 2 Regular graphs
Consider a finite field $\mathbb{F}_{q}$ with $q$ ($q$ prime) elements, and the vector space $\mathbb{F}_{q}^{k}$ over $\mathbb{F}_{q}$. Choose randomly $d/2$ invertible matrices $T_{1},\ldots,T_{d/2}$, and use the corresponding permutations of the set of non-zero tuples to get an undirected graph of degree $d$: a non-zero tuple $x\in\mathbb{F}_{q}^{k}$ has neighbors $T_{1}x,\ldots,T_{d/2}x,T_{1}^{-1}x,\ldots,T_{d/2}^{-1}x.$ This is a regular graph with $q^{k}-1$ vertices and degree $d$ (that may contain loops and parallel edges), and we compute numerically the second eigenvalue of this graph.
This procedure is repeated many times, and then the empirical distribution of the second eigenvalues is shown (Figure 1). For comparison, we show also the distribution for the random permutation model and for a different field sizes (and approximately the same graph size).
Figure 1: All the eigenvalues (horizontal axis) are divided by $2\sqrt{d-1}$. The two upper curves (quite close to each other) show the empirical distributions for random permutation model (dashed line) and random linear permutation model for graphs of degree $d=30$ ($15$ permutations) and size $116383=2^{14}-$ 1. The third curve (wider green one in the bottom part) shows the distribution for random linear permutations on $5$-tuples from a $7$-element field, so the graph has $168067^{5}-1=$ 16806 vertices (and the same degree $30$, obtained with $15$ permutations). For each distribution $5000$ experiments were made; the eigenvalues are grouped in $40$ bins.
The invertible matrices were generated by choosing a random [^1] The random number generator was Mersenne twister ([17]); we tested physical random generators without noticing any difference. matrix and then testing whether it is invertible. For Boolean matrices, the fraction of invertible ones is above $1/4$ (see, e.g., [1]), so we do not need too many trials; for bigger fields, the fraction is even bigger. Computing the determinant (checking the invertibility) is easy. For computing the eigenvalues we used the C++ library Spectra [^2] Spectra’s home page: https://spectralib.org/. that implements the Lanczos algorithm [7].
@@
We also tested another set of parameters: $5$-tuples from a $7$-element field (so the resulting graph size is close to the previously considered one, $168067^{5}-1=$ 16806 instead of $163832^{14}-1=$ 16383, which makes the results roughly comparable). Note that this choice allows us to decrease the number of random bits used for the graph generation. We see that this is achieved for a price: the distribution is now less concentrated, though still very close to the Ramanujan threshold. (One could expect that more random bits lead to better concentration.) One can also observe (we do not have any explanation for this paradoxical observation) that the probability to get a random graph that is on the left of, say, $0.998$, becomes bigger for a bigger field. So, if we try to get one graph with small eigenvalue after many trials, the bigger field may be better. One could try to go to the extreme in this direction and consider $1\times 1$ invertible matrices, i.e., consider multiplication by non-zero elements in some field $\mathbb{F}_{q}$ (the graph size is then $q-1$). This is, however, a bad idea since in this case the invertible matrices commute and therefore many paths in a random walk arrive at the same place just due to commutativity. Such a graph will have poor mixing properties.
Additionally, we can observe that variations on the degree of the graphs (for a given construction) does not seem to have much impact on the distribution of the second largest eigenvalues (Figure 2, left). Though the difference is hardly visible, we can see that smaller degree leads to a slightly greater second largest eigenvalue. This is not surprising since low degree random graphs tend to be worse expanders. However, the eigenvalues are slightly more concentrated. As expected (see Figure 2 on the right) increasing the dimension does not change the expected value but decreases the variance.
Figure 2: On the left: second largest eigenvalue distributions for $\textit{GL}_{14}(\mathbb{F}_{2})$ Schreier graphs of degrees $10,30,60$ and $120$. On the right: instead of the degree, we vary the dimension of the matrices and keep the same field $\mathbb{F}_{2}$.
@@ 3 Bipartite biregular graphs
The experiments in the previous section were for undirected regular graphs, but one may also need a bipartite graph that is biregular (degrees of the vertices are the same in each part, but may differ in left and right parts) and has good mixing properties (for the back and forth random walk).
@@
However, in the bipartite case we do not need to use a permutation together with its inverse. Instead, we consider $d$ random bijections between parts and get a biregular graph of degree $d$. We use this construction in our experiments with bipartite graphs (the previous construction would have given the same result as for regular non-bipartite graphs). This allows us to have bipartite graphs of any (possible odd) degree.
@@
Figure 3: Here, n1 is the size of each partition. The dashed line (blue) shows the distribution of second eigenvalues for a bipartite graph constructed from $30$ random permutations of $16383$-element set (nonzero bit vectors of size $14$). The orange line (close to the first one) shows the same for $30$ linear permutations (over $\mathbb{F}_{2}$). The third line (green, in the bottom part) shows the distribution for $30$ permutations of a $16806$-element set that are randomly chosen among the $\mathbb{F}_{7}$-linear permutations of nonzero $5$-tuples with elements in $\mathbb{F}_{7}$.
These constructions can be performed both with random permutations and random linear permutations. The same question arises: what can be said about mixing properties of graphs in these two settings? A traditional way to measure mixing property of a bipartite graph is to consider the (non-normalized) adjacency matrix for it (height and width are equal to the total number of vertices in both parts). The eigenvalues of this matrix are grouped into pairs, starting with $\sqrt{d_{L}d_{R}}$, where $d_{L}$ and $d_{R}$ are left and right degrees (the eigenvector coordinates are $\sqrt{d_{L}}$ in the left part and $\sqrt{d_{R}}$ in the right part) and $-\sqrt{d_{L}d_{R}}$ (to change the sign of the eigenvalue we change signs in all eigenvector coordinates in one of the parts). The next pair, $\lambda>0$ and $-\lambda<0$, determines the mixing properties, and we call $\lambda$ the second largest eigenvalue (ignoring the negative ones).
@@
For the case of different left and right degrees the results of our numerical experiments are shown in Figure 4. Comparing this figure with the previous one, we observe some differences: the random permutations now give slightly (but visibly) better results compared to random linear permutations. On the other hand, changing the size of the field (and matrix dimension) now has much less effect. In fact, results may depend on which triples we are merging, for the linear case, due to the additional structure induced by the the construction (compared to general permutations). However, this phenomenon still has to be explained.